# Pathology-CoT: Learning Visual Chain-of-Thought Agent from Expert Whole Slide Image Diagnosis Behavior

### v1

@@ Abstract
Diagnosing a whole-slide image diagnosis is an interactive, multi-stage process of changing magnification and moving between fields. Although recent pathology foundation models are strong, practical agentic systems that decide what field to examine next, adjust magnification, and deliver explainable diagnoses are still lacking. The blocker is data: scalable, clinically aligned supervision of expert viewing behavior that is tacit and experience‑based, not written in textbooks or online, and therefore absent from LLM training. We introduce the AI Session Recorder, which works with standard WSI viewers to unobtrusively record routine navigation and convert the viewer logs into standardized behavioral commands (inspect/peek at discrete magnifications) and bounding boxes. A lightweight human-in-the-loop review turns AI-drafted rationales into the Pathology-CoT dataset, a form of paired “where to look” and “why it matters” supervision produced at a roughly six-fold lower labeling time. Using this behavioral data, we build Pathologist-o3, a two-stage agent that first proposes ROIs and then performs behavior-guided reasoning. On gastrointestinal lymph-node metastasis detection it achieved 84.5% precision, 100.0% recall, and 75.4% accuracy, exceeding the state-of-the-art OpenAI o3 model and generalizing across backbones. To our knowledge, this constitutes one of the first behavior-grounded agentic systems in pathology. Turning everyday viewer logs into scalable, expert‑validated supervision, our framework makes agentic pathology practical and establishes a path to human‑aligned, upgradeable clinical AI.
@@ 1 Introduction
The advancement of digital pathology and its integration of artificial intelligence (AI) has deeply intertwined over the past two decades [1, 2]. This progress enabled the first generation of computer-aided diagnosis tools, some of which are now incorporated into clinical workflows for tasks such as tumor detection and cell counting [3, 4, 5, 6]. These early successes demonstrated the value of AI in the treatment of well-defined problems within pathology. More recently, the field has been transformed by the emergence of powerful self-supervised pathology foundation models [7, 8, 9] alongside large language models (LLMs). Together, these models combine visual understanding with language-level reasoning, opening the way to the next frontier: AI pathology agents.
LLM-based agents have already begun to transform complex biomedical domains such as genetic experiment design and clinical diagnosis [10, 11, 12, 13]. In whole-slide image (WSI) analysis, an agentic paradigm is particularly well suited, since diagnosis is a dynamic, multi-step process rather than a static classification task. Here, an agent refers to a system capable of perceiving its environment, leveraging LLM to perform reasoning and experiment orchestration, and executing a sequence of actions, such as zooming, panning, and selecting regions, to accomplish a complex diagnostic goal. Consequently, an agent that can systematically examine and analyze gigapixel WSIs provides a natural foundation for a collaborative AI partner for pathologists.
Although designing an agent system plays a role, recent advances such as OpenAI o3 [14] and its variants [15, 16] have already established capable agent frameworks. In digital pathology, to ensure alignment with pathologists’ routine operations and achieving scalability, the main bottleneck is the availability of high-quality, well-curated training data. Unlike programming or mathematical agents, clinical diagnosis cannot be guided by simple rule-based rewards. Agents instead need to learn how pathologists examine slides and which regions they attend to for each task [17, 18]. This underscores the urgent need to encode human knowledge and practice into reusable computational forms, which in turn requires direct expert involvement in creating such training data. However, gigapixel WSIs encompass diverse diagnostic scenarios and workflows [19, 20], often makes dataset construction especially time-consuming, costly, and difficult to scale.
Modern digital pathology systems routinely record navigation events such as panning and zooming, thereby creating a continuous record of pathologist behavior [20]. Although these interaction logs are produced at scale, such user behavior signals are both high-frequency and noisy, and thus require careful post-processing before they can serve as effective supervisory signals [21, 22] (Figure 1a). These “digital exhaust” traces can provide a scalable source of supervision for training human-aligned agents, but only if they are accurately captured, curated, and converted into structured, model-ready representations.
To address this critical gap, we introduce the AI Session Recorder, which converts raw digital pathology interaction logs into a structured, agent system-ready pathology-CoT dataset. The proposed system segments continuous event streams and discretizes pathologists’ interactions into a compact sequence of actions. These actions include exploratory scans at low or medium magnification and brief high-magnification inspections, each linked to LLM-friendly coordinates that define a Region of Interest (ROI). For each ROI, a vision–language model (VLM) will generate a concise rationale and impression, which the pathologist can quickly verify or edit through a human-in-the-loop interface (Figure 1b). This process extracts signals from noise, producing data that are well suited for training LLM-based agents. By coupling the AI Session Recorder with pathologist oversight, annotation costs are substantially reduced while maintaining clinical fidelity. In the gastrointestinal lymph node metastasis task of the Pathology-CoT dataset, this paradigm achieved an eight-fold increase in the efficiency of labeling.
To evaluate the effectiveness of behavioral data, we developed pathologist-o3, a task-conditioned agent trained on recorded viewing behavior. The model achieved 84.5% precision, 100.0% recall, and 75.4% accuracyaaaThe lower accuracy compared to both precision and recall is due to the dataset’s high imbalance., outperforming the state-of-the-art closed-source commercial thinking-with-image OpenAI o3 model, which reached 46.7% precision, 87.5% recall, and 57.8% accuracy. Across different backbones, incorporating behavioral guidance increased both precision and recall by an average of 11.8% and 17.9%, respectively. In addition, the agent demonstrated greater navigation completeness (61.3%) and efficiency (56.2%), much higher when compared with baseline, unguided VLMs. Our approach also differs fundamentally from PathChat+ [23], which delegates the selection of where to zoom and why to an LLM "supervisor." The procedural viewing policy that pathologists use is largely tacit and not captured in text corpora, making it difficult for text‑trained LLMs to acquire. Consistent with this limitation, PathChat+’s navigation actions generated by LLM that diverge from expert trajectories and often lack diagnostic utility (Fig. 4). By directly learning from recorded expert navigation and pairing it with pathologist‑verified rationales, we provide explicit supervision for both the "where" and the "why," enabling human‑aligned behavior and stronger end‑to‑end performance.
We make three key contributions throughout this paper. First, we introduce digital pathology AI Session Recorder, which captures routine viewer navigation and, with human-in-the-loop review, transforms raw logs into standardized, agent-ready training data. Second, we introduce the pathology-CoT dataset. It provides task-conditioned, expert-validated “behavior + reasoning” pairs, delivering clinically aligned supervision at scale. Third, we develop the pathologist-o3 agent, it can identify diagnostic regions and reasons over standardized crops, outperforming unguided VLM baselines and transferring successfully to an external validation cohort.
With these contributions, we tackle the central bottleneck for innovating and advancing digital pathology agents: obtaining data that is both clinically meaningful and scalable. By converting routine viewing logs into task-conditioned behavioral supervision paired with low-cost, expert-verified reasoning, we supply the supervision required for existing architectures to evolve into capable, human-aligned agents.
@@
Figure 1: The AI Session Recorder framework for converting expert viewing behavior into agent-ready training data. a The AI Session Recorder addresses the core data challenge: raw, continuous user logs are too noisy and complex for AI to learn from. The recorder’s algorithm transforms this chaotic data stream into discrete, meaningful commands (e.g., ‘5x-Inspect’), inspired by the discrete objective lenses (e.g., 5x, 10x, 40x) on physical microscopes. b The recorder’s novel data generation pipeline creates the Pathology-CoT dataset. For each captured action, an AI drafts a rationale explaining the expert’s focus, which a pathologist then efficiently verifies or corrects in a human-in-the-loop workflow. c The resulting dataset contains 5,222 conversational rounds from pathologists with diverse experience levels. d The semi-automated, human-in-the-loop workflow is highly efficient, reducing labeling time by approximately six-fold compared to manual annotation from scratch.
@@ 2.1 A scalable data engine converts viewing behavior into agent-ready supervision
Diagnosing a whole-slide image for conditions like lymph node metastasis is a dynamic process of visual inquiry. Pathologists begin with a low-power overview, zoom into suspicious areas for high-power analysis, and iteratively navigate across the gigapixel canvas to build a conclusion (Figure 1a left). This sequence of interactions is highly purposeful; the keyboard and mouse movements embody the pathologist’s tacit knowledge, reflecting an expert’s experience-driven strategy for searching for evidence. While modern digital pathology systems routinely capture this behavior in interaction logs, the raw data exists as a high-frequency, noisy, and enormously long stream. This “digital exhaust” is too complex and voluminous for direct use in training machine learning models, as the chaotic input exceeds the context limits of current AI and prevents effective learning. This gap means that the vast expertise exercised in daily clinical practice remains in a format that is not amenable to model training. A systematic methodology to record, structure, and encode this knowledge is fundamentally required to train sophisticated AI agents that can serve as true diagnostic co-pilots or autopilots.
@@
Given this urgent challenge, we developed the pathology “AI session recorder”, a tool that aims at resolving the data bottleneck of training next-generation medical AI agent by collecting pathologists behavior data at scale. AI session recorder processes the behavior data from 8 pathologists at Stanford Medicine when using nuclei.io [24] to make diagnosis for adjacent lymph node metastasis. Such raw data contains 25 cases, 137 slides across 8 pathologists (four attending pathologists, two fellows, and two residents) with a total of 10.6 hours of inspection. Based on such data, AI session recorder performs two key functions to transform unusable digital pathology image inspection raw logs into structured, agent-ready supervision (Figure 1a). First, it automatically analyzes the complex viewing process to identify and discretize the most important areas of focus (Figure 1b). Second, inspired by the discrete objective lenses of a physical microscope, this process converts the chaotic event stream into a compact sequence of meaningful behavioral commands. Each command is defined as either a broad, low-magnification <inspect> or a rapid, high-magnification <peek>, and is paired with a standardized region-of-interest (ROI) box. The recorder also captures clinical reasoning by using these expert‑identified ROIs to prompt a vision‑language model (VLM), which drafts a rationale explaining why the region was examined and what key findings are present (Figure 1b, left). The model‑generated rationales are reviewed and, when necessary, edited by two pathologists, then consolidated into the Pathology‑CoT agent‑training dataset (Figure 1b, right, shows an example).
@@
To implement and verify this AI session recorder, we focused on a clinically representative task, namely N-staging of colorectal cancer (CRC) lymph-node metastasis, which is both common and labor-intensive. We collected behavioral data from eight pathologists at Stanford Medicine as they analyzed 137 WSIs using a standard viewer that recorded all navigation events [24]. The raw logs were noisy and high-frequency, with active navigation captured at approximately 10 Hz and averaging more than 257 distinct viewport events per slide (Figure S1). Using this stream directly as supervision was infeasible because feeding the raw viewports into a VLM would produce over 500,000 visual tokens per case, far exceeding the current context limits and could also prevent effective learning. In contrast, the AI session recorder reduces excessive context length by filtering noisy, non-informative behavior data, while preserving the fidelity of pathologists’ behavioral signals. Using this tool, human pathologists can rapidly verify AI-drafted behavior and rationale by accepting, revising, or rejecting them.
@@
Based on the manual filtering, the resulting Pathology-CoT dataset is a structured collection of expert-validated, task-conditioned “behavior + reasoning” pairs. Derived from 10.6 hours of recorded interactions from eight pathologists with diverse experience levels (attendings, fellows, and residents) across 921 sessions (Figure 1d), the dataset contains 5,222 conversational rounds of agent training data. The reasoning is detailed: broad, low-magnification <inspect> actions are accompanied by an average of 152 words of description, while the analysis of fine cellular-level findings from high-magnification <peek> actions averages 82 words. This rich dataset captures not only where and how an expert looks but also embeds the clinical reasoning for why, providing a robust, multi-modal foundation for training and evaluating pathology agents. To our knowledge, this is the first chain-of-thought pathology behavior data that captured and established from expert pathologists. Different from PathChat+ [23] that the knowledge was based on GPT model, the thinking ability is based on expert pathologists.
Figure 2: Overview of pathology-o3 and its performance. a Workflow: a task prompt initializes the agent, which proposes candidate regions using a behavior bank/locating module; cropped ROIs are sent to a VLM that generates per‑ROI reasoning, then summarized into an impression and final diagnosis. b Example “thinking with image” on one slide. The agent lists planned inspections, provides step‑wise descriptions for each ROI (blue boxes), and synthesizes a case‑level conclusion. Positive and negative findings are indicated for each step. c Quantitative comparison on the lymph‑node metastasis task. Pathologist-o3 achieves the best overall accuracy with perfect or near‑perfect recall, outperforming strong VLM baselines (asterisks denote significance).
@@ 2.2 Behavior-guided supervision enables accurate pathology agents
AI session recorder enables a scalable way to record tacit diagnostic knowledge from pathologists at scale. By that we are able to collect a scalable dataset “Pathology-CoT” that is ready for machine learning. Based on such format of dataset, it allows (i) digitize and store pathologists’ behavior; (ii) guiding the action model (such as LLM) to locate important regions from WSI; and (iii) training an agentic AI system to make clinical diagnosis in the same manner as pathologists.
To demonstrate the utility of that, we designed a task-conditioned agent “pathology-o3” that mimics a pathologist’s workflow. Learned from the training data “Pathology-CoT dataset” in example of lymph node metastasis, the agent first decides where to look, then analyzes what it sees (Figure 2a). To emulate this first step, it uses a task-specific Behavior Predictor trained on the complete set of discretized behavioral commands from aforementioned Pathology-CoT dataset, then learning to propose diagnostic ROIs on a new slide in our internal and external validation datasets. Once these ROIs are identified, the system performs behavior-guided thinking: for each region, it auto-crops a standardized high-resolution field, injects task context into the prompt (case-level goal and information, and ROI-level intent), and then prompts a VLM to generate per-ROI reasoning, which is subsequently synthesized into a case-level summary. This capability by combining a learned human viewing policy with a powerful reasoning AI, allows the agent to produce a step-by-step visual chain-of-thought for a given whole-slide, demonstrating its ability to systematically analyze a slide in a human-aligned manner (Figure 2b).
We evaluated our framework against a panel of state-of-the-art VLMs, with detailed specifications provided in Table S1. On the lymph-node metastasis benchmark, our behavior-guided agent achieved 84.5% precision, 100% recall, and 75.4% accuracy (Figure 2c). The next-best model, OpenAI o3, reached 46.7% precision, 87.5% recall, and 57.8% accuracy. Text-centric VLMs performed notably worse and exhibited imbalanced trade-offs: for example, GPT-4.1 achieved high recall (91.7%) but low precision (42.3%), while Llama-4 showed higher precision (75.0%) but extremely low recall (6.2%). Across the remaining baselines, precision ranged from 33.3–65.6%, recall from 18.2–91.7%, and accuracy from 42.2–65.6%. These results demonstrate the importance of grounding reasoning in expert visual search patterns; for nuanced diagnostic challenges like lymph node metastasis, generic reasoning is insufficient. Performance hinges on first learning where and how to look (i.e., thinking with image), which our behavior-aligned data supplies.

### v2

@@ Abstract
Diagnosing a whole-slide image is an interactive, multi-stage process of changing magnification and moving between fields. Although recent pathology foundation models demonstrated superior performances, practical agentic systems that decide what field to examine next, adjust magnification, and deliver explainable diagnoses are still lacking. Such limitation is largely bottlenecked by data – scalable, clinically aligned supervision of expert viewing behavior that is tacit and experience‑based, not documented in textbooks or internet, and therefore absent from LLM training. Here we introduce Pathology-CoT, a framework designed to address this challenge through three key breakthroughs. First, the AI Session Recorder seamlessly integrates with standard whole-slide image (WSI) viewers to unobtrusively record routine navigation and convert the viewer logs into standardized behavioral commands (inspect/peek at discrete magnifications) and bounding boxes. Second, a lightweight human-in-the-loop review turns AI-drafted rationales for behavioral commands into the Pathology-CoT dataset, a form of paired “where to look” and “why it matters”, enabling six-fold faster labeling compared to manual constructing such Chain-of-Thought dataset. Using this behavioral data, we build Pathology-o3, a two-stage agent that first proposes important regions of interest and then performs behavior-guided reasoning. On the gastrointestinal lymph-node metastasis detection task, our method achieved 100.0% recall on the internal validation dataset from Stanford Medicine and 97.6% recall on an independent external validation dataset from Sweden, exceeding the state-of-the-art OpenAI o3 model and generalizing across backbones. To our knowledge, Pathology-CoT constitutes one of the first behavior-grounded agentic systems in pathology. Turning everyday viewer logs into scalable, expert‑validated supervision, our framework makes agentic pathology practical and establishes a path to human‑aligned, upgradeable clinical AI.
@@ 1 Introduction
The advancement of digital pathology and its integration of artificial intelligence (AI) have become deeply intertwined over the past two decades [1, 2]. This progress enabled the first generation of computer-aided diagnosis tools, some of which are now incorporated into clinical workflows for tasks such as tumor detection and cell counting [3, 4, 5, 6]. These early successes demonstrated the value of AI in some of well-defined problems within pathology. More recently, the field has been transformed by the emergence of powerful self-supervised pathology foundation models [7, 8, 9] alongside large language models (LLMs) and vision language models (VLMs). Together, these models combine visual understanding with language-level reasoning, opening the way to the next frontier: AI pathology agents.
@@
Figure 1: The AI Session Recorder framework for converting expert viewing behavior into agent-ready training data. a The AI Session Recorder addresses the core data challenge: raw, continuous user logs are too noisy and complex for AI to learn from. The recorder’s algorithm transforms this chaotic data stream into discrete, meaningful commands (e.g., 5x-Inspect), inspired by the discrete objective lenses (e.g., 5x, 10x, 40x) on physical microscopes. b The recorder’s novel data generation pipeline creates the Pathology-CoT dataset. For each captured action, an AI drafts a rationale explaining the expert’s focus, which a pathologist then efficiently verifies or corrects in a human-in-the-loop workflow. c The Pathology-CoT dataset contains 5,222 conversational rounds from pathologists with diverse experience levels. d The semi-automated, human-in-the-loop workflow is highly efficient, reducing labeling time by approximately six-fold compared to manual annotation from scratch.
LLM-based agents have already begun to transform complex biomedical domains such as genetic experiment design and clinical diagnosis [10, 11, 12, 13]. In whole-slide image (WSI) analysis, an agentic paradigm is particularly well suited, since diagnosis is a dynamic, multi-step process rather than a static classification task. Here, an agent refers to a system capable of perceiving its environment, leveraging an LLM to perform reasoning and experiment orchestration, and executing a sequence of actions, such as zooming, panning, and selecting regions, to accomplish a complex diagnostic goal. Consequently, an agent that can systematically examine and analyze gigapixel WSIs provides a natural foundation for a collaborative AI partner for pathologists.
Although such model have a complex architecture, recent advances such as OpenAI o3 [14] and its variants [15, 16] have already established capable agent frameworks with image zoom-in Chain-of-Thought (CoT) capability. In digital pathology, to ensure alignment with pathologists’ routine operations and achieving scalability, the main bottleneck is the availability of high-quality, well-curated training data. Unlike programming or mathematical agents, clinical diagnosis cannot be guided by simple rule-based rewards. Agents instead need to learn how pathologists examine slides and which regions they attend to for each task [17, 18]. While several datasets are now available, consisting of cropped image-caption pairs crawled from the internet [19] or pairings of clinical WSIs and reports [20], these sources typically lack pathologist viewing behavior. Such lack of information underscores the urgent need to encode human knowledge and practice into reusable computational forms, which in turn requires direct expert involvement in creating such training data. However, gigapixel WSIs encompass diverse diagnostic scenarios and workflows [21, 22], often make dataset construction especially time-consuming, costly, and difficult to scale.
Modern digital pathology systems routinely record navigation events such as panning and zooming, thereby creating a continuous record of pathologist behavior [22]. Although these interaction logs can be produced at scale, such user behavior signals are both high-frequency and noisy, and thus require careful post-processing before they can serve as effective supervisory signals [23, 24] (Figure 1a). These “digital exhaust” traces can provide a scalable source of supervision for training human-aligned agents, but only if they are accurately captured, curated, and converted into structured, model-ready representations.
To address this gap, we introduce the Pathology-CoT framework, which comprises three key innovations. The AI Session Recorder segments continuous event streams and discretizes pathologists’ behavior into a compact sequence of conversational text. These actions include exploratory scans at low or medium magnification and brief high-magnification inspections, each linked to LLM-friendly coordinates that define a Region of Interest (ROI). For each ROI, a VLM will generate a concise rationale and impression, which the pathologist can quickly verify or edit through a human-in-the-loop interface (Figure 1b). The resulting outputs collectively form the Pathology-CoT dataset, a high-quality behavioral corpus that captures not only where experts look but also why specific regions matter. This process distills meaningful behavioral signals from raw viewing logs, producing data ideally suited for training LLM-based agents. By integrating automated recording with expert validation, the AI Session Recorder substantially reduces annotation costs while maintaining clinical fidelity. In the gastrointestinal lymph node metastasis task within the Pathology-CoT dataset, this paradigm achieved a six-fold improvement in labeling efficiency. To evaluate the effectiveness of behavioral data, we developed Pathology-o3, an agent trained on recorded viewing behavior in Pathology-CoT. The model achieved 84.5% precision, 100.0% recall, and 75.4% accuracy, outperforming the state-of-the-art closed-source commercial thinking-with-image OpenAI o3 model, which reached 46.7% precision, 87.5% recall, and 57.8% accuracy. Across different backbones, incorporating behavioral guidance increased both precision and recall by an average of 11.8% and 17.9%, respectively. Furthermore, the agent’s learned viewing policy demonstrated strong generalization, maintaining robust performance (97.6% recall) on a challenging independent external validation cohort. Critically, Pathology-o3’s navigation strategy was highly aligned with that of a senior attending pathologist—both qualitatively and quantitatively—a stark contrast to the scattered, low-yield navigation of other VLMs.
With these contributions, this study tackle the central bottleneck for innovating and advancing digital pathology agents: obtaining data that is both clinically meaningful and scalable. By converting routine viewing logs into task-conditioned behavioral supervision paired with low-cost, expert-verified reasoning, we supply the supervision required for existing architectures to evolve into capable, human-aligned agents. The Pathology-CoT framework thus represents a pivotal step toward behavior-grounded, clinically trustworthy AI systems in computational pathology.
@@ 2.1 A scalable data engine converts viewing behavior into agent-ready supervision
Diagnosing a whole-slide image for conditions like lymph node metastasis is a dynamic process of visual inquiry. Pathologists begin with a low-power overview, zoom into suspicious areas for high-power analysis, and iteratively navigate across the gigapixel canvas to build a conclusion (Figure 1a left). This sequence of interactions is highly goal-oriented: the keyboard and mouse movements embody the pathologist’s tacit knowledge, reflecting an expert’s experience-driven strategy for searching for evidence. While modern digital pathology systems routinely capture this behavior in interaction logs, the raw data exists as a high-frequency, noisy, and enormously long stream. Such “digital exhaust” is too complex and massive for direct use in training machine learning models, as the chaotic input exceeds the context limits of current AI and prevents effective learning. This gap means that the vast expertise exercised in daily clinical practice remains in a format that is not amenable to model training. A systematic methodology to record, structure, and encode this knowledge is fundamentally required to train sophisticated AI agents that can serve as true diagnostic co-pilots or autopilots.
@@
Given this urgent challenge, we developed the pathology “AI Session Recorder”, a tool that aims at resolving the data bottleneck of training next-generation medical AI agent by collecting pathologist’s behavior data at scale. AI Session Recorder performs two key functions to transform unusable digital pathology image inspection raw logs into structured, agent-ready supervision (Figure 1a). First, it automatically analyzes the complex viewing process to identify and discretize the most important areas of focus (Figure 1b). Second, inspired by the discrete objective lenses of a physical microscope, this process converts the chaotic event stream into a compact sequence of meaningful behavioral commands. Each command is defined as either a broad, low-magnification <inspect> or a rapid, high-magnification <peek>, and is paired with a standardized region-of-interest (ROI) box. The recorder also captures clinical reasoning by using these expert‑identified ROIs to prompt a vision‑language model (VLM), which drafts a rationale explaining why the region was examined and what key findings are present (Figure 1b, left). The model‑generated rationales are reviewed and, when necessary, edited by two pathologists, then consolidated into the Pathology‑CoT agent‑training dataset (Figure 1b, right, shows an example).
@@
To implement and verify this AI Session Recorder, we focused on a clinically representative task, namely N-staging of colorectal cancer (CRC) lymph-node metastasis, which is both common and labor-intensive. We collected behavioral data from eight pathologists at Stanford Medicine using a standard viewer nuclei.io [25] that recorded all navigation events of making diagnosis for adjacent lymph node metastasis in colorectal cancer. Such raw data contains 25 cases, 137 slides across 8 pathologists (four attending pathologists, two fellows, and two residents) with a total of 10.6 hours of inspection (Figure 1c). The raw logs were noisy and high-frequency, with active navigation captured at approximately 10 Hz and averaging more than 257 distinct viewport events per slide (Figure S1). Using this stream directly as supervision was infeasible because feeding the raw viewports into a VLM would produce over 500K visual tokens per case, far exceeding the current context limits and could also prevent effective learning. In contrast, the AI Session Recorder reduces excessive context length by filtering noisy, non-informative behavior data, while preserving the fidelity of pathologists’ behavioral signals. AI Session Recorder is also robust, showing similar behavior pattern across same pathologist’s two viewing session (Figure S2). Using this tool, human pathologists can rapidly verify AI-drafted behavior and rationale by accepting, revising, or rejecting them.
@@
Based on the manual filtering, the resulting Pathology-CoT dataset is a structured collection of expert-validated, task-conditioned “behavior + reasoning” pairs. Pathology-CoT dataset is a large-scale public dataset derived from 10.6 hours of recorded interactions from eight pathologists with diverse experience levels (attendings, fellows, and residents) across 921 sessions (Figure 1c), the dataset contains 5,222 conversational rounds of agentic vision CoT training data. The reasoning is detailed: broad, low-magnification <inspect> actions are accompanied by an average of 152 words of description, while the analysis of fine cellular-level findings from high-magnification <peek> actions averages 82 words. This rich dataset captures not only where and how an expert looks but also embeds the clinical reasoning for why, providing a robust, multi-modal foundation for training and evaluating pathology agents. Unlike PathChat+ [26], whose reasoning was based on VLM-generated text, our dataset captures the thinking process of expert pathologists. To our knowledge, this is the first chain-of-thought pathology behavior data that captured and established from expert pathologists.
Figure 2: Overview of Pathology-o3 and its performance. a Workflow: a task prompt initializes the agent, which proposes candidate regions using a behavior bank/locating module; cropped ROIs are sent to a VLM that generates per‑ROI reasoning, then summarized into an impression and final diagnosis. b Example “thinking with image” on one slide. The agent lists planned inspections, provides step‑wise descriptions for each ROI (blue boxes), and synthesizes a case‑level conclusion. Positive and negative findings are indicated for each step. c Quantitative comparison on the lymph‑node metastasis task. Pathology-o3 achieves the best overall accuracy with perfect or near‑perfect recall, outperforming strong VLM baselines. Error bars represent 95% bootstrap confidence intervals.
@@ 2.2 Behavior-guided supervision enables accurate pathology agents
AI Session Recorder enables a scalable way to record tacit diagnostic knowledge from pathologists at scale. By that we are able to collect a scalable “Pathology-CoT dataset” that is ready for machine learning. Based on such format of dataset, it allows (i) digitize and store pathologists’ behavior; (ii) guiding machine learning model to locate important regions from a WSI; and (iii) training an agentic AI system to make clinical diagnosis in the same manner as pathologists.
To demonstrate the utility of that, we designed a task-conditioned agent “Pathology-o3” that mimics a pathologist’s workflow. Learned from the training data Pathology-CoT in example of lymph node metastasis, the agent first decides where to look, then analyzes what it sees (Figure 2a). To emulate this first step, it uses a task-specific Behavior Predictor trained on the complete set of discretized behavioral commands from aforementioned Pathology-CoT dataset, then learning to propose diagnostic ROIs on a new slide in our internal and external validation datasets. Once these ROIs are identified, the system performs behavior-guided thinking: for each region, it auto-crops a high-resolution field, injects task context into the prompt (case-level goal and information, and ROI-level intent), and then prompts a VLM to generate per-ROI reasoning, which is subsequently synthesized into a case-level summary. This capability by combining a learned human viewing policy with a powerful reasoning AI, allows the agent to produce a step-by-step visual chain-of-thought for a given WSI, demonstrating its ability to systematically analyze a slide in a human-aligned, easy to understand and validate manner (Figure 2b).
We evaluated our framework against a panel of state-of-the-art VLMs. On the lymph-node metastasis benchmark, our behavior-guided agent achieved 84.5% precision, 100% recall, and 75.4% accuracy (Figure 2c). The next-best model, OpenAI o3, reached 46.7% precision, 87.5% recall, and 57.8% accuracy. VLMs without chain-of-thought ability performed notably worse and exhibited imbalanced trade-offs: for example, Phi-4 achieved high recall (91.7%) but low precision (42.3%), while Llama-4 showed higher precision (75.0%) but extremely low recall (6.2%). Across the remaining baselines, precision ranged from 33.3–65.6%, recall from 18.2–91.7%, and accuracy from 42.2–65.6%. These results demonstrate the importance of grounding reasoning in expert visual search patterns; for nuanced diagnostic challenges like lymph node metastasis, generic reasoning is insufficient. Performance hinges on first learning where and how to look (i.e., thinking with image), which our behavior-aligned data supplies. More detail on the setting for VLM and performance metrics are shown in Table S1 and Table S2. It is important to note that multiple instance learning (MIL) and related approaches require scanning every patch of a whole-slide image, making the process computationally intensive and fundamentally different from the goal-oriented behavior of human experts. Because such exhaustive patch-level processing cannot be efficiently guided or learned from human behavioral data, these MIL-based methods were not included in the comparison.